OpenStudy (anonymous):

Find the modulo class to which the number belongs for the given modulo system. 69, mod 15 A. 2 B. 13 C. 9 D. 12

jimthompson5910 (jim_thompson5910):

Find the remainder when you divide 69 into 15

jimthompson5910 (jim_thompson5910):

What is 69 divided by 15?

OpenStudy (anonymous):

i got 4.6

jimthompson5910 (jim_thompson5910):

Alright, so 69 goes into 15 four whole times This means that 69/15 = 4*15 + r where r is some whole number and it is the remainder

jimthompson5910 (jim_thompson5910):

what is the value of r?

OpenStudy (anonymous):

I dont understand where r is coming from ?

jimthompson5910 (jim_thompson5910):

r is that leftover remainder when the number doesn't fully go into 15

jimthompson5910 (jim_thompson5910):

If the number did evenly go into 15, then that remainder would be zero (but it's not)

OpenStudy (anonymous):

I got 3.75

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Imagine 15 as the lenth of a circle, then wrap a rope of length 69 around it. It would wrap around 4 complete times but wouldn't make it around the 5th time. the left over portion would represent the modulo...how much would you have left over?

OpenStudy (anonymous):

Think back to elementary school division. If you were dividing 7 by 3, you wouldnt put 2.3333 You would say 2 with a remainder of 1, Since 3 goes into 7 twice, but 2*3 = 6, and it takes one more to get to 7. Here are some more examples: 16/7 = 2 with a remainder of 2, since 2*7 is 14, and it takes 2 more to get to 16 19/8 = 2 with a remainder of 3 since 2*8 = 16, and it takes 3 more to get to 19. We are asking, what is the remainder when you do 69/15? that will be your answer.